OpenStudy (anonymous):

Find the slope, if it exists, of the line containing the pair of points. (-16, -2) and (-20, -18)

OpenStudy (anonymous):

It looks like the slope exists since the difference of x-values does not equal to zero.

OpenStudy (rosedewittbukater):

m = y2 - y1/x2 - x1 m = -18-(-2)/-20-(-16) m = -18+2/-20+16 m = -16/-4 m = 4 The slope is 4

OpenStudy (anonymous):

To find the slope when two points are given, use the Slope Formula. Check out this example from my website. http://chilimath.com/algebra/intermediate/line/slope-formula.html#example1

OpenStudy (anonymous):

the slope, "m", can be found using the slope formula:\[m = \frac{ y _{1} - y _{2} }{ x _{1} - x _{2} }\]where you use your given points that are in the form:\[(x _{1}, y _{1})\]and\[(x _{2}, y _{2})\]

OpenStudy (anonymous):

So, x1=-16, y1=-2, x2=-20, and y2=-18 So, you substitute those values into the formula. Anytime that the 2 x's are not the same, you will get a slope. The slope will be undefined when the x's are the same and in that case, the line will be vertical.

OpenStudy (anonymous):

Thank you all so much

OpenStudy (anonymous):

You're quite welcome!

OpenStudy (anonymous):

but then it says to type an interger or simplified fraction

OpenStudy (anonymous):

The formula to find the slope is m = (y2-y1)/(x2-x1). Substituting the values we get... \[m=\frac{ -18-(-2) }{ -20-(-16) }\] \[=\frac{ -16 }{ -4 }\] \[=4\]

OpenStudy (anonymous):

Once you make the subsitutions, you will get either an integer or a fraction that may or may not need simplification.

OpenStudy (anonymous):

Making sense now?

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

That's the important part, when a person sees how its done and can do other problems from it. That's a good formula to memorize. Good luck in all of your studies!

OpenStudy (anonymous):

Thank you, I have more to post but I am going to close this question